RD-2018-03-26

From Electron Ion Collider
Jump to: navigation, search

EIC Tracking R&D Meetings: Mondays 9:00am - 11:00pm

To join the meeting on a computer or mobile phone: https://bluejeans.com/703875605

Want to test your video connection? http://bluejeans.com/111

--

Connecting directly from a room system?

1) Dial: 199.48.152.152 or bjn.vc

2) Enter Meeting ID: 703875605

--

Just want to dial in on your phone?

1) +1.408.740.7256 (US)

+1.888.240.2560 (US Toll Free)

+1.408.317.9253 (Alternate number)

(http://bluejeans.com/numbers)

2) Enter Meeting ID: 703875605 3) Press #


Link to the previous meeting: https://wiki.bnl.gov/eic/index.php/RD-2018-03-12

Link to all previous meetings: https://wiki.bnl.gov/eic/index.php/Eic-rd-meeting


Kondo, Marcus, Matt


Progress at BNL

  • ...

Progress at FIT (Marcus)

  • Test the gas leak issue on the mock up chamber with the new longer screws. Long screws help a little bit but does not eliminate the gas leak but so far no gas at the output.
  • Marcus suspects that major leak is happening at the corners of the CF frames. Will try stronger screws for tighter closing of the chamber
  • Also looking at adding some washers to the screw all along the frames to improve sealing. The group is working on all these possible solutions
  • Marcus gives a tip on how to recover a stubborn GEM foil HV sector that is failing the HV at 500V. One can try to quickly apply 1kV across the GEM but in a dry environment by flowing CO2 or N2 in the chamber

Progress at INFN Trieste

  • ...

Progress at SBU (Klaus by e-mail)

  • ... Progress status regarding mirror coating is the same: it is in the making.

Progress at UVa (Kondo)

  • EIC weeklyMeeting20180326.pdf will be presented in two weeks from now
  • Report on the effort on the support frames. Parts of low cost frames production done by machine shop and gluing of all 4 frames was done in the detector lab
  • High cost frames are under production at RESARM. Expect to be delivered in about two weeks
  • Drawing of the modification of the PRad stretcher to be re-used for EIC prototype was completed and submitted to the machine shop. Expected to have them in about 2 weeks from now.
  • Marcus asks for the UG student for the simulation. Replied that we are still working on it

Progress at Temple U. (Matt)

  • Still waiting for Amilkar returns from duties at BNL to start the assembly of a 3rd prototype
  • Kondo asks about the current status of the second prototype with straight spacers. Matt replied that they lost 2 HV sectors and the chamber is no longer taking data, HV is down. They are investigating the problem
  • Matt is scanning old Tech Etch GEM foils (spare foils from the BoNUS experiment at JLab) provided by Howard Fenker. The BoNus RTPC using these GEMs and experiences huge gain non uniformity. foil size are ~ 20cm x 15cm. Howard wanted to check whether this was due to nun uniform GEM holes geometry. Matt measurement of one foil shows pretty good uniform holes distribution with a width similar to what is observed with a typical CERN GEM foil.
  • Matt suspect that a non uniform gain might be due to the Kapton material used at that time by Tech Etch rather than the Apical used by CERN and now by Tech Etch. Kondo suspect that it has more to do with how the BoNUs chamber was built (gas circulation, uniformity of the gap between GEM in a cylindrical geometry) rather than the material.
  • Marcus ask Matt if he would be willing to test Korean large GEM foil now produced using the double mask technique. This is OK with Matt

AOB

  • ...